OpenStudy (anonymous):

Proof that a^2>0=>a>0.

OpenStudy (mertsj):

How can you prove something that isn't true?

OpenStudy (chriss):

right... any number squared is greater than 0. example let a=-2 -2^2=4 so a<0 and a^2>0

OpenStudy (callisto):

for any real number a^2>0 but \[\sqrt{a ^{2}}\] can be a or -a

OpenStudy (callisto):

for complex number \[i ^{2}=-1\] where \[i=\sqrt{-1}\]

OpenStudy (callisto):

anyway that's not true

OpenStudy (mertsj):

Is this horse dead yet?

OpenStudy (anonymous):

Thanks for all.

OpenStudy (mertsj):

yw